OpenStudy (anonymous):

Show that -1,3,6 are the zeroes of the polynomial f(x)=x^3-8x^2+9x+18. Also verify the relationship b/w the zeroes and coefficient of f(x)

OpenStudy (anonymous):

HELPMEPLEASEEEEEEEEEEEEEEEEEEEE

OpenStudy (anonymous):

What you have to do is set up your formula to equal 0, then plug in -1,3,6 for X.

OpenStudy (tkhunny):

Have you considered substituting the provided values? Synthetic division would help speed that up. Have you considered factoring the polynomial. That might be the fastest of all.

OpenStudy (wolf1728):

Using -1 -1 ^3 = -1 -8*(-1^2) = -8 9*x = -9 +18 = 18 Sums to 0 3 and 6 will also sum to zero.